What is "x" in 8/9x+5=-3
Answer by checkley79(3341)

8/9x+5=-3
8X/9=-3-5
8X/9=-8 CROSS MULTIPLY.
8X=-8*9
8X=-72
X=-72/8
X=-9 ANS.